How many electrons leave the battery per minute?
Answer #1

V=9 V

Resistance R= 2.7 ohm.

Curve (i)=V/R=9/2.7A .............(1)

Current is charge leaving the battery per sec.

i=Q/t

Q=it ............(2)

Magnitude of char on electron =e=1.6x10-19 c.

Let number of electrons leaving the battery is n.

Charge on n electrons:-

Q=ne=it ( from equation 2 )

t=one minute =60 sec

n=it/e =(9/2.7)x60/(1.6x10-19 ) =200/(1.6x10-19)

n=1.25x1021 electrons Ans.